if di + va = 0 (0 represents the zero vector) and || 01 || = || 0 || = 273, calculate v.uz
Answer #1

业 业 业 业 Given that +22=0 and will=11V2ll2253. since J1 +22=0. → Y + 31 -0. 22. J + Z27=0 (by inner product). <,+》十x VL,方+ 20.